OpenStudy (anonymous):

Determine whether the following aare subspaces of C[-1,1]: (a) The set of functions f in C[-1,1] such that f(-1)=f(1) What is the procedure followed to proove these kind of questions ?

OpenStudy (zarkon):

I assume you know the definition of a subspace?

OpenStudy (anonymous):

yes

OpenStudy (zarkon):

what does it tell you ... that you need to show?

OpenStudy (anonymous):

closure under vector addition, scalar mult and that there is the 0 vector in it

OpenStudy (zarkon):

ok

OpenStudy (zarkon):

so if \(f\in C[-1,1]\) and \(f(-1)=f(-1)\) does it follow that \[cf\in C[-1,1]\]

OpenStudy (zarkon):

\(f(-1)=f(1)\)

OpenStudy (zarkon):

ie is \(cf(-1)=cf(1)\)

OpenStudy (anonymous):

i don't really understand why they need to point out that f(-1) = f(1)

OpenStudy (zarkon):

they are just creating a subspace with certain properties.

OpenStudy (zarkon):

if the axioms hold then it is a subspace

OpenStudy (zarkon):

so is \(cf(-1)=cf(1)\)?

OpenStudy (anonymous):

true

OpenStudy (zarkon):

if \(f,g\in C[-1,1]\) does it hold that \(f+g\in C[-1,1]\)?

OpenStudy (anonymous):

not necessarly

OpenStudy (zarkon):

we should also note that if \(f,g\) are continuous then so are \(cf\) and \(f+g\)

OpenStudy (zarkon):

why?

OpenStudy (anonymous):

oh ok, i understand what you meant

OpenStudy (zarkon):

if \(f,g\in C[-1,1]\) then \(f(-1)=f(1)\) and \(g(-1)=g(1)\) thus \(f(-1)+g(-1)=f(1)+g(1)\)

OpenStudy (zarkon):

ok?

OpenStudy (anonymous):

I understand that

OpenStudy (zarkon):

what is the zero vector?

OpenStudy (anonymous):

f(0) ?

OpenStudy (zarkon):

no...\(f(x):=0\)

OpenStudy (zarkon):

then \(f\) is continuous and \(f(-1)=0=f(1)\) thus \(f\in C[-1,1]\)

OpenStudy (anonymous):

thank, for the help but who did u know that f(−1)=0=f(1) ?

OpenStudy (zarkon):

if \(f\) is the zero vector...ie the zero function then \(f(x)=0\) for all \(x\). That includes -1 and 1